OpenStudy (kainui):

A slightly interesting and simple integral, can anyone tell me why this is true?

OpenStudy (kainui):

\[\int\limits_{}^{} \tan (x) \sec^2 (x) dx \] So if you pick u=sec(x) you get a different answer than if you pick u=tanx. Why is that?

OpenStudy (dls):

The change is adjusted by the integration constant C which you apply normally I guess,if you take some interval and evaluate it as a definite integral you will get the same answer no matter what term you let=u.

OpenStudy (kainui):

Yeah, I guess I was looking more for the fact that: tan^2x+1=sec^2x so taking the derivative of both sides will give you the same thing.

ganeshie8 (ganeshie8):

1/2sec^2x + c1 = 1/2tan^2x + c2

ganeshie8 (ganeshie8):

what DLS said is also same...

ganeshie8 (ganeshie8):

1 is absorbed by C

OpenStudy (kainui):

Out of curiosity when you see that integral what does everyone see first the tangent or secant integral?

ganeshie8 (ganeshie8):

myself u = tanx ofcourse :) u = secx dint occur to me till date... it kept me thinkning few seconds.. when u asked to use u = secx

OpenStudy (anonymous):

Yes DLS is right if you do the first method you get answer as \[I = \frac{(tanx)^2}{2} = \frac{\sec^2x}{2}-\frac{1}{2}\]

OpenStudy (kainui):

Yeah, I'm just looking for clever/interesting integrals. It is sort of interesting to think that tan^2x has the same derivative as sec^2x, I suppose when i think of them they seem like they should be much more different from each other, but really one is just shifted up by 1 of the other haha.

ganeshie8 (ganeshie8):

lol u gona become an evil professor someday :P

OpenStudy (kainui):

Good. Err... Bad. =)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=tan%5E2%28x%29 http://www.wolframalpha.com/input/?i=sec%5E2%28x%29 but they have the same graph :D.. no wonder they have the same slope.. !!

OpenStudy (kainui):

I'm noticing something interesting here, this implies an infinite definition...\[\sec^2x=1+\sin^2x \sec^2x\]\[\sec^2x=1+\sin^2x (1+\sin^2x (1+\sin^2x (...)))\] I don't know how interesting or useful that is, just something to think about I guess haha.

OpenStudy (anonymous):

\[1+\sin^2x(1+\sin^2x(....) = y \\1+(\sin^2x)(y) =y \\ y-ysin^2x= 1 \\ycos^2x =1\\ y= \sec^2x\] just for the heck of it xD

OpenStudy (kainui):

Haha I just finished at the same time as you posted that myself. Fun stuff, actually you know the trig identities for sin^2x, cos^2x, and sinxcosx in terms of a single a trig function? I found a fairly intuitive way of deriving them if you're interested since I hate memorizing them but now I can fairly easily just come up with them in my head lol.

OpenStudy (anonymous):

m gonna go now :P.. have fun with math xD

OpenStudy (anonymous):

here is something interesting do u know what is 1+2+3+4..... upto infinity?

OpenStudy (anonymous):

The answer is -1/12 :D.. true story!

OpenStudy (kainui):

Yeah, but we all know geometric series diverges outside of -1 and 1. ;P

OpenStudy (kainui):

S=1+x+x^2+x^3+... xS=x+x^2+x^3+... 1+xS=1+x+x^2+x^3+...=S 1+xS=S 1=S(1-x) S=1/(1-x) =P